OpenStudy (anonymous):

Solve for the variable: x/3 - 1/2 = 5/6 STEPS PLEASE

OpenStudy (anonymous):

x/3=5/6+1/2 x/3= 5/6 + 3/6 x/3=8/6 x=8/2 x=4

OpenStudy (anonymous):

In step 2 why did you change 1/2 to 3/6?

OpenStudy (anonymous):

so tht 5/6 and 1/2 can add together

OpenStudy (anonymous):

ohhhh I get ya.